Provide the structure of the major fragment that results when the molecular ion of
CH3CH2CH2CH2OH undergoes fragmentation via alpha cleavage.
Answer #1

CH3-CH2-CH2 - CH2 – OH Il d-clevage CH2 - CH2 - CH – CH275 m 12 = 74 H OH t-H2O CHE -GH2 – CH=CH2 7+ [m-18] M/2 = 56 molecula